I have 7 quizzes each quiz is worth 100 points the average that I need is a 75 what is the lowest grade I can make on each quiz to get an average of 75 points ?
Answer by TimothyLamb(4379)

x = sum of scores for 7 quizzes
x/7 = 75
x = 75*7
x = 525
---
525/100 = 5.25
---
answer:
there are many ways to average 75 over 7 quizzes, these are two ways:
1) get 100 on 5 quizzes, 25 on one quiz, and zero on another quiz.
2) get 75 on all 7 quizzes.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php